OpenStudy (anonymous):

Find the limit. lim x→−∞ square root of 4x^6 − x divided by x^3 + 4

OpenStudy (freckles):

\[\lim_{x \rightarrow - \infty}\frac{\sqrt{4x^6-x}}{x^3+4}\]?

OpenStudy (freckles):

but actually what I translate you said is this: \[\lim_{- \rightarrow - \infty}(\sqrt{4x^6}-\frac{x}{x^3}+4 )\]

OpenStudy (freckles):

but for some reason I don't think it is that

OpenStudy (xapproachesinfinity):

if it is the first one @freckles you can pull out x^6 on the nominator and x^3 in the denominator

OpenStudy (freckles):

you mean divide both top and bottom by sqrt(x^6} but sqrt(x^6)=-x^3 so this needs to be kept in mind

OpenStudy (xapproachesinfinity):

you can also pull out as i said! and yes you have to keep in mind that we taking negative values of x